OpenStudy (anonymous):

Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $61 to produce 3 calculators, $165 to produce 7 calculators, and $385 to produce 12 calculators. Find the total cost of producing 6 calculators

OpenStudy (anonymous):

please help

OpenStudy (amistre64):

you are given 3 points to match a quadratic equation to. you will simply need to develop 3 equations in 3 unknowns and solve it by whatever method you like best

OpenStudy (amistre64):

take the general form: ax^2 + bx + c = y, and plug in the 'points' given a(3)^2 + b(3) + c = 61 a(7)^2 + b(7) + c = 165 a(12)^2 + b(12) + c = 385 substitution, elimination, or matrix are the 3 methods most commonly employed